# Personalized Tree based progressive regression model for watch-time prediction in short video recommendation

### v1

@@ Abstract.
The tree structure serves as the core of TPM, as it defines the decomposition of watch time estimation and determines how the ordinal intervals are discretized. However, in TPM, the tree is predefined as a full binary tree, which may be sub-optimal for the following reasons. First, a full binary tree implies an equal partitioning of the watch time space, which may struggle to capture the complexity of real-world watch time distributions. Second, instead of relying on a globally fixed tree structure, we advocate for a personalized, data-driven tree that can be learned in an end-to-end manner. Therefore, we propose PTPM to enable a highly personalized decomposition of watch estimation with better efficacy and efficiency. Moreover, we reveal that TPM is affected by selection bias due to conditional modeling and devise a simple approach to address it. We conduct extensive experiments on both offline datasets and online environments. The results of the offline experiment illustrate the improvement in watch time accuracy, and the results of the online AB test further validate the effectiveness of the proposed framework. PTPM has been fully deployed in our core traffic scenarios and currently serves more than 400 million users per day.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In video-sharing platforms (e.g., TikTok and Kuaishou), watch time denotes the amount of time users spend on watching videos, which is strongly related to watch time and Daily Active Users (DAU). Watch time prediction is thus an essential task in online video recommender systems.
@@
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Watch time normalized by durationQuantileMusicGameMakeupOutfitFunSportsBeauty
@@
  • PTPM supports real-time generation of personalized trees for individual requests, while achieving a computational efficiency similar to that of standard TPM (see Fig. 6 and Table 4). This is particularly important in industrial recommender systems with strict latency requirements.
@@ 2. Related Work
Ordinal Regression Ordinal regression addresses problems where target labels have a natural order, such as age estimation (Chang et al., 2011) and search relevance prediction (Wu et al., 2003), and is often implemented by converting ordinal targets into binary classification tasks. Following this idea, CREAD (Sun et al., 2024) defines each classifier to indicate whether the label exceeds a threshold. TPM (Lin et al., 2023) organizes ordinal intervals into a tree structure, where watch time prediction is formulated as traversing root-to-leaf paths. This hierarchical design not only preserves the ordinal nature of watch time, but also models conditional dependencies and uncertainty in a structured way.
A major challenge in ordinal regression is determining appropriate thresholds, an issue that remains underexplored. Among existing studies, only CREAD (Sun et al., 2024) explicitly addresses the discretization of continuous watch time. It relies on a hand-crafted heuristic formula to determine discretization thresholds and tunes hyperparameters via grid search over equal-width and equal-frequency schemes. However, CREAD suffers from several limitations, including fixed and non-adaptive discretization, a constrained heuristic search space, and the high computational cost of grid search due to repeated training. In contrast, our proposed PTPM jointly trains the tree structure and its associated classifiers to learn a context-aware partitioning strategy.
Tree Structure Learning Traditional tree-based models like XGBoost (Chen and Guestrin, 2016) use greedy algorithms to construct trees that partition the feature space. In contrast, our method focuses on dividing the task space for adaptive watch time decomposition, which is a fundamental difference. JTM (Zhu et al., 2019), an extension of TDM (Zhu et al., 2018), jointly learns the tree structure and user representations. PTPM differs in three key ways: (1) It uses the tree for task decomposition and watch time estimation, while JTM focuses on item retrieval; (2) It enables end-to-end tree learning, avoiding the greedy, layer-wise construction in JTM; (3) It generates input-specific trees on the fly, unlike JTM’s static, shared structure.
@@ 3.1. Watch Time Prediction Task
$N_{T}:\{n_{i}\}$ set of nodes in ${T}$
$L_{T}:\{l_{k}\}$ set of leaf nodes in ${T}$
@@ 3.2. The TPM Approach
Uncertainty modeling. TPM not only predicts the expected watch time but also quantifies the uncertainty of its predictions. Since watch time is treated as a random variable following a categorical distribution $p(\hat{y}_{i}\in l_{k}\mid x_{i},T)$ for all $l_{k}\in L_{{T}}$, the uncertainty is measured by the variance:
@@ 4.1.1. Formulation
[…]
@@
[…]
@@ 4.1.2. Optimizing $\mathcal{L}_{tree}$
where $R_{T}$ is the evaluation metric of the tree $T$. We simply set $R_{T}=\text{XAUC}(y,\hat{y})-\text{MSE}(y,\hat{y})$ (see Sec. 5.1.3).
@@ 4.2. Unbiased Conditional Learning
Formally, the ideal cross entropy loss $\delta$ for a classifier is: $\mathcal{R}=\frac{1}{|\mathcal{X}|}\sum_{x_{i}\in\mathcal{X}}\delta(o_{i},\hat{o_{i}})$, where $\mathcal{X}$ denotes the fully observed user-item pairs, $o_{i}\in\{0,1\}$ is the label $\mathbb{I}(y_{i}\in{n}_{\phi_{l_{k}}}(j)|x_{i})$ and $\hat{o_{i}}=p(\hat{y}_{i}\in{n}_{\phi_{l_{k}}}(j)|x_{i})$ is the prediction. However, in TPM, to capture the conditional prior, $\hat{o}_{i}$ can only be trained for samples that $y_{i}\in{n}_{\phi_{l_{k}}}(j-1)$:
(13) ^$\hat{\mathcal{R}}_{native}=\frac{1}{|x_{i}:y_{i}\in{n}_{\phi_{l_{k}}}(j-1)|}\sum_{x_{i}:y_{i}\in{n}_{\phi_{l_{k}}}(j-1)}\delta(o_{i},\hat{o_{i}})$
We call this the native estimator because it introduces selection bias, making ^$\hat{\mathcal{R}}_{native}$ a biased estimate of the true performance $\mathcal{R}$, i.e. $\mathbb{E}_{\mathbb{I}}[\hat{\mathcal{R}}_{native}]\neq\mathcal{R}$, where $\mathbb{I}$ is short for $\mathbb{I}(y_{i}\in{n}_{\phi_{l_{k}}}(j))$.
@@
(15) ^$\hat{\mathcal{R}}_{IPS}=\frac{1}{|\mathcal{X}|}\sum_{x_{i}:y_{i}\in{n}_{\phi_{l_{k}}}(j-1)}\frac{\delta(o_{i},\hat{o_{i}})}{p(\hat{y}_{i}\in n_{\phi_{l_{k}}}(j-1)|{x}_{i})}$
@@ 5.2.1. Comparison with other methods
#Parameters#Computation(flops)20202020404040406060606032.2732.2732.2732.2764.5564.5564.5564.5532.5532.5532.5532.5565.165.165.165.1Value (M)TPMPTPM
@@
11114444777710101010131313131616161619191919222222222525252528282828313131310.50.50.50.511111.51.51.51.52222Leaf node $l_{k}$ in $T_{g}$ with depth 6Ratio of prior-to-posteriorIdeal: $y=1.0$TPMTPM+UCL
TPMTPM+UCL0.550.550.550.550.60.60.60.60.650.650.650.650.590.590.590.590.620.620.620.62Avg. AUC of Classifiers
@@
33334444555566660.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.65Depth $d(T_{g})$XAUCPTPMPTPM w.t.o. TSL
33334444555566663333444455556666Depth $d(T_{g})$Avg. depth of PTPM

### v2

# Personalized Tree-Based Progressive Regression Model for Watch-Time Prediction in Short Video Recommendation

@@ Abstract.
The tree structure is central to TPM, as it defines the decomposition of watch time estimation and how ordinal intervals are discretized. However, TPM uses a predefined full binary tree, which may be sub-optimal for two reasons. First, full binary trees imply equal partitioning of the watch time space, which may fail to capture the complexity of real-world distributions. Second, rather than relying on a fixed global structure, we advocate for a personalized, data-driven tree that can be learned end-to-end. Thus, we propose PTPM to enable highly personalized decomposition of watch estimation with better efficacy and efficiency. Moreover, we show that TPM suffers from selection bias due to conditional modeling and propose a simple solution. We conduct extensive experiments on offline datasets and online environments. Offline results show improved watch time accuracy, and online A/B tests further validate the effectiveness of our framework. PTPM has been fully deployed in core traffic scenarios and now serves over 400 million users daily.
Watch time prediction; Recommendation; Tree models
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761538isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommender systems
@@ 1. Introduction
In video-sharing platforms (e.g., TikTok and Kuaishou), watch time denotes the amount of time users spend on watching videos, which is strongly related to user engagement and Daily Active Users (DAU). Watch time prediction is thus an essential task in online video recommender systems.
@@
00.20.20.40.40.60.60.80.81100.20.20.40.40.60.60.80.811Watch time normalized by durationQuantileMusicGameMakeupOutfitFunSportsBeauty
@@
  • PTPM supports real-time personalized tree generation per request, while maintaining computational efficiency comparable to standard TPM. This is crucial for industrial recommender systems with strict latency requirements.
@@ 2. Related Work
Ordinal Regression Ordinal regression addresses problems where target labels have a natural order, such as age estimation (Chang et al., 2011) and search relevance prediction (Wu et al., 2003), and is often implemented by converting ordinal targets into binary classification tasks. Following this idea, CREAD (Sun et al., 2024) defines each classifier to indicate whether the label exceeds a threshold. TPM (Lin et al., 2023) organizes ordinal intervals into a tree structure, where watch time prediction is formulated as traversing root-to-leaf paths.
A major challenge in ordinal regression is determining appropriate thresholds, an issue that remains underexplored. Among existing studies, only CREAD (Sun et al., 2024) explicitly addresses the discretization of continuous watch time. It uses a hand-crafted heuristic formula to determine thresholds and tunes hyperparameters via grid search over equal-width and equal-frequency schemes. However, CREAD has several limitations, including fixed, non-adaptive discretization, a constrained heuristic space, and the high computational cost of grid search due to repeated training. In contrast, our proposed PTPM jointly trains the tree structure and its associated classifiers to learn a context-aware partitioning strategy.
Tree Structure Learning Traditional tree-based models like XGBoost (Chen and Guestrin, 2016) use greedy algorithms to construct trees that partition the feature space. In contrast, our method divides the task space for adaptive watch time decomposition—a fundamental difference. JTM (Zhu et al., 2019), an extension of TDM (Zhu et al., 2018), jointly learns the tree structure and user representations. PTPM differs in three key ways: (1) It uses the tree for task decomposition and watch time estimation, while JTM focuses on item retrieval; (2) It enables end-to-end tree learning, avoiding JTM’s greedy, layer-wise construction; (3) It generates input-specific trees on the fly, unlike JTM’s static, shared structure.
@@ 3.1. Watch Time Prediction Task
$N_{T}:\{n_{i}\}$ set of nodes in T{T}
$L_{T}:\{l_{k}\}$ set of leaf nodes in T{T}
@@ 3.2. The TPM Approach
Uncertainty modeling. TPM predicts not only expected watch time but also the uncertainty of its predictions. Treating watch time as a random variable following a categorical distribution $p(\hat{y}_{i}\in l_{k}\mid x_{i},T)$ for all $l_{k}\in L_{{T}}$, uncertainty is measured by the variance:
@@ 4.1.1. Formulation
(5) $\text{Outer:}\quad T^{*}=\arg\min_{T}\mathcal{L}_{\text{tree}}(\mathbf{\theta}^{*}(T),T)$
$\text{Inner:}\quad\mathbf{\theta}^{*}(T)=\arg\min_{\boldsymbol{\theta}}\mathcal{L}_{\text{TPM}}(\boldsymbol{\theta},T)$
@@
(a) $\text{s.t.}\quad{\mathbf{\theta}}^{*}(T)$ $={\mathbf{\theta}}^{*}(T_{g})\odot(1-\mathbf{a}_{T})$
@@ 4.1.2. Optimizing $\mathcal{L}_{tree}$
[…]
@@ 4.2. Unbiased Conditional Learning
[…]
We call this the native estimator because it introduces selection bias, making $\hat{\mathcal{R}}_{native}$ a biased estimate of the true performance $\mathcal{R}$, i.e. $\mathbb{E}_{\mathbb{I}}[\hat{\mathcal{R}}_{native}]\neq\mathcal{R}$, where $\mathbb{I}$ is short for $\mathbb{I}(y_{i}\in{n}_{\phi_{l_{k}}}(j))$.
@@
(15) $\hat{\mathcal{R}}_{IPS}=\frac{1}{|\mathcal{X}|}\sum_{x_{i}:y_{i}\in{n}_{\phi_{l_{k}}}(j-1)}\frac{\delta(o_{i},\hat{o_{i}})}{p(\hat{y}_{i}\in n_{\phi_{l_{k}}}(j-1)|{x}_{i})}$
@@ 5.2.1. Comparison with other methods
#Parameters#Computation(flops)20204040606032.2732.2764.5564.5532.5532.5565.165.1Value (M)TPMPTPM
@@
114477101013131616191922222525282831310.50.5111.51.522Leaf node $l_{k}$ in $T_{g}$ with depth 6Ratio of prior-to-posteriorIdeal: y=1.0y=1.0TPMTPM+UCL
TPMTPM+UCL0.550.550.60.60.650.650.590.590.620.62Avg. AUC of Classifiers
@@
334455660.50.50.550.550.60.60.650.65Depth $d(T_{g})$XAUCPTPMPTPM w.t.o. TSL
3344556633445566Depth $d(T_{g})$Avg. depth of PTPM
@@ 6. Conclusion
7. GenAI Usage Disclosure
No generative AI tools were used in any part of this work, including the design of experiments, code implementation, data analysis, and manuscript writing.